OpenStudy (anonymous):

5. What is the vertex in the equation x^2 – 6x + 4 explain

OpenStudy (anonymous):

if the vertex is the point where the graph intersect the x axis, then set the equation equals to zero and solve for x

OpenStudy (helder_edwin):

complete the square.

OpenStudy (anonymous):

find the minimum of the function; that would be ur vertex, sorry dont set it to 0

OpenStudy (helder_edwin):

no. complete the square. do u know how?

OpenStudy (anonymous):

take the first derivative... dont complete the sqaure

OpenStudy (anonymous):

2x-6 =0 so x=3 plug 3 in the original eqtion, u get -5 (3,-5) is your vertex

OpenStudy (helder_edwin):

ok. u have \[ \large y=x^2-6x+4 \] \[ \large y-4+\color{red}{3^2}=x^2-6x+\color{red}{3^2} \] \[ \large y+5=(x-3)^2 \]

OpenStudy (helder_edwin):

so the vertex is (3,-5)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.algebra-class.com/vertex-formula.html

OpenStudy (anonymous):

so if you have ax^2+bx+c ,based on the sign of a, the parabola opens up or down. if a is positive, the parabola opens up; if a is negative it opens down... so here a is 1>0 so it opens up

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

\[x1 = (-b+\sqrt{b^2 - 4*a*c})/(2*a); x2=(-b-\sqrt{b^2 - 4*a*c})/(2*a)\]

OpenStudy (anonymous):

plugging in the values of both a, b and c, you will get your x1 and x2

OpenStudy (anonymous):

good. now can you take it from there?

OpenStudy (anonymous):

sorry but im kinda confused with the way you wrote your answer. hold a second

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

using a calculator unless you can simplify it which is what I am doing now

OpenStudy (anonymous):

x1=(6+√6^2−(4∗1∗4))/(2∗1) x2=(6−√6^2−(4∗1∗4))/(2∗1)

OpenStudy (anonymous):

do you understand that?

OpenStudy (anonymous):

thx